Given that f(2)=4,f(2.5)=5.5 obtain the linear interpolation polynomial using Newton's Divided Difference


1
Expert's answer
2021-05-11T14:35:58-0400

Newton's divided difference table is:




Newton's divided difference interpolation formula is:

"f(x)=y_0+(x-x_0)f[x_0,x_1]\\f(x)=4+(x-2)\times 3\\f(x)=4+(3x-6)\\f(x)=3x-2"


Answer:

Polynomial: y = 3x-2